candiceowens87:

A sporting goods store sells sets of golf clubs with steel shafts for $309. Sets of golf clubs with graphite shafts cost $489. Last month the store sold 11 sets of golf clubs for $4,299. How many sets of each type of golf club were sold? Set up and solve a system of equations to find the answer.

justus:

Hello there. Let's put s for steal and g is for graphite So we have s+g=11 and 309s+489g=4299 Now we solve for g g=11-s then this gives us 309s+489(11-s)=4299 309s+5379-489s=4299 -180s+5379=4299 -180s=-1080 s=6 g=5

justus:

Also s is 6 since s+g+11